# A Survey of Self-Evolving Agents: On Path to Artificial Super Intelligence

### v3

@@ 5.3.2 Multi-Agent Evolution
System Architecture Evolution. This paradigm focuses on evolving the structural and coordination aspects of multi-agent systems, including team composition, orchestration strategies, and workflow optimization. EvoMAC hu2024self introduces a framework that mimics neural network training for multi-agent systems, implementing "textual backpropagation" where compilation errors and test failures serve as loss signals to drive iterative modifications of agent team composition and individual prompts. A specialized "updating team" analyzes textual feedback to identify problematic agents and generate modification instructions, effectively implementing gradient-based optimization in the space of agent configurations rather than model parameters. Building on this structural evolution concept, Puppeteer dang2025multi takes a different approach by focusing on coordination strategy evolution rather than team composition changes. The system employs a centralized orchestrator that evolves its decision policy through reinforcement learning, dynamically selecting which agents to activate at each step while balancing task performance with computational cost. This "puppeteer-puppet" paradigm demonstrates how architectural evolution can occur at the coordination level, discovering efficient collaboration patterns and emergent behaviors such as tighter coordination among core agents and sophisticated cyclic interaction patterns.
@@
Knowledge-Based Evolution. This paradigm emphasizes evolving the collective knowledge and experience of multi-agent teams through memory accumulation and case-based learning, primarily operating through in-context learning or in-context-like adaptation rather than parameter updates. MDTeamGPT chen2025self establishes the foundation for this approach through a dual knowledge base system, implementing CorrectKB for storing successful cases and ChainKB for capturing failure reflections, enabling the system to learn from both successes and mistakes through structured case retrieval and reasoning enhancement. Extending this medical consultation framework, MedAgentSim almansoori2025self demonstrates how such knowledge-based evolution can be applied to real-world diagnostic scenarios, accumulating experience from patient interactions and using retrieval-augmented generation to improve consultation quality over time. PiFlow pu2025piflow applies this paradigm to scientific discovery, maintaining a trajectory of principle-outcome pairs and using them to steer hypothesis generation through information-theoretical optimization.
@@ 5.4 Cross-cutting Evolutionary Dimensions
Figure 7: Illustration of cross-cutting evolutionary dimensions in agent self-evolution, structured along three key axes: learning paradigm (offline/online), policy consistency (on/off-policy), and reward granularity (process-based, outcome-based, and hybrid). These dimensions jointly characterize how autonomous agents generate data, interact with environments, adapt policies, and receive feedback, providing a structured lens for analyzing reward-based, imitation-based, and population-based evolution strategies.
@@ 5.4.1 Online and Offline Learning
Offline Learning In the offline learning paradigm, the learning phase is decoupled from live task execution. The offline process typically involves cycles of offline data generation, filtering, and model fine-tuning, focusing on building a powerful and generalist foundational model before deployment. A primary strategy in this domain is LLM bootstrapping, where a model enhances its own capabilities using its self-generated content. For example, Self-Instruct wang2022self shows how a language model can bootstrap its own instruction-following ability by generating new instructions, paired with its own responses, creating a synthetic dataset for fine-tuning. Building on this, WizardLM xu2024wizardlm demonstrates how to progressively evolve the complexity of these self-generated instructions, pushing the model’s capabilities on more challenging tasks. In the context of GUI and Web agents, offline learning often involves leveraging pre-collected high-quality trajectories for supervised fine-tuning (SFT). OS-Genesis sun2024genesis introduced a reverse task synthesis method for automatic trajectory creation. Similarly, UI-Genie xiao2025ui employs a unified reward model for trajectory evaluation and a self-improving loop to generate high-quality trajectories iteratively. Both approaches focus on curating a rich SFT dataset to enhance the agent’s capabilities to solve complex tasks. Beyond SFT, offline methods also incorporate reinforcement learning performed on a static dataset of agent-environment interactions. For example, GUI-R1 luo2025gui and InfiGUI-R1 liu2025infigui utilize rule-based rewards and apply R1-style guo2025deepseek training on offline GUI datasets.
@@
Online Learning In contrast, online learning enables an agent to learn and adapt continuously while it interacts with a live or simulated environment. Feedback from each action is used to update the agent’s policy, plan, or knowledge base in real-time. This allows for greater adaptability to dynamic or unseen situations. Some agents evolve online not by updating their model weights, but by refining their plans and skill libraries on the fly. For example, Voyager wang2023voyager presents an LLM-powered agent that learns to play Minecraft by continuously exploring, generating its own curriculum of tasks, and building a persistent skill library from direct experience. AdaPlanner sun2023adaplanner focuses on adapting its plan within a task; it generates an initial plan, receives feedback from the environment, and refines the plan online. Similarly, SwiftSage lin2023swiftsage operates with a fast-and-slow thinking process, where it can reflect on failures of its fast, intuitive mode and switch to a more deliberate, tool-using slow mode, adapting its strategy online based on task difficulty. Reinforcement Learning serves as a fundamental mechanism for online learning, enabling agents to learn from environmental reward signals. DigiRL bai2024digirl demonstrates how to train device-control agents in the wild using autonomous RL, while DistRL wang2024distrl proposes an asynchronous distributed framework to make such on-device training feasible. MobileGUI-RL shi2025mobileguirladvancingmobilegui addresses the specific challenges of training GUI agents in online mobile environments by introducing a synthetic task generation pipeline combined with group relative policy optimization (GRPO) through trajectory-aware rewards.
@@ 5.4.2 On-policy and Off-policy Learning
On-policy Learning. On-policy approaches require agents to learn exclusively from experiences generated by their current policy, ensuring policy consistency but often at the cost of sample efficiency. Reflexion shinn2023reflexion exemplifies this approach through its iterative self-reflection mechanism. The agent generates responses using its current policy, receives feedback on failures, and immediately incorporates this feedback to update its reasoning process for the next iteration. GRPO shao2024deepseekmath and DAPO yu2025dapo continue this path and show the effectiveness of multiple rollouts. The agent always learns from its current behavior, maintaining strict policy consistency. In agent settings, on-policy methods provide excellent learning stability and avoid distribution mismatch issues that plague off-policy methods. However, they suffer from low sample efficiency, as each policy update requires fresh data collection, making them computationally expensive for complex multi-step reasoning or tool use scenarios where generating high-quality trajectories is costly.
Off-policy Learning. Off-policy approaches allow agents to learn from experiences generated by different policies, including previous versions, other agents, or human demonstrations, significantly improving sample efficiency at the cost of potential distribution mismatch. yuan2401self demonstrates a sophisticated off-policy approach where model $M_{t+1}$ learns from preference data generated by the previous version $M_{t}$. The system handles distribution shift through DPO’s built-in KL divergence constraint with the reference policy, preventing the new policy from deviating too far from the data-generating policy. yuan2023rrhf showcases another powerful off-policy paradigm by learning from diverse response sources—including other models, humans, and different sampling strategies—through ranking-based supervision. The method elegantly sidesteps distribution shift by treating alignment as a ranking problem rather than requiring policy consistency. zhao2025sirius illustrates off-policy learning in multi-agent settings, where agents learn from an "experience library" containing successful interaction trajectories generated by previous policy versions, enabling efficient reuse of expensive multi-agent coordination data. In agent settings, off-policy methods excel in sample efficiency, allowing agents to leverage historical data, expert demonstrations, and cross-agent learning. They are particularly valuable for multi-step reasoning where successful trajectories are rare and expensive to generate, and for tool use scenarios where agents can learn from diverse execution examples without repeated environmental interaction. However, they face challenges with distribution shift, reward hacking (where agents exploit inconsistencies between training and deployment policies), and the need for careful regularization to maintain training stability.
@@ 5.4.3 Reward Granularity
Outcome-based Reward Outcome-based Reward is a feedback mechanism that evaluates an agent based on the successful completion of predefined tasks. This reward is determined solely by the final state of the agent’s trajectory, regardless of the intermediate steps. A central challenge, particularly in dynamic environments like web or GUI navigation, is to effectively learn from both successful trajectories and the much more frequent failure trajectories. To address this, Direct Preference Optimization (DPO) rafailov2023direct is designed to directly maximize the likelihood of preferred responses while minimizing the KL-divergence with the reference policy. Similarly, RRHF yuan2023rrhf employs a ranking loss approach that aligns model probabilities of multiple responses with human preferences by ranking response probabilities without requiring auxiliary value models. Moreover, several works have developed specialized frameworks for agent self-evolution that are built upon outcome-based rewards. A straightforward approach is rejection sampling finetuning, as used in AutoWebGLM lai2024autowebglm . This method employs a pre-designed reward model to evaluate trajectory outcomes, identify the successful trajectories, and update the model with this high-quality data. DigiRL bai2024digirl models the GUI navigation task as a Markov Decision Process (MDP) and obtains a final, sparse reward at the end of an episode using a VLM-based evaluator. WebRL qi2024webrl develops a robust outcome-supervised reward model (ORM) to address the feedback sparsity inherent in dynamic web environments. The ORM evaluates task success within a self-evolving curriculum framework, enabling agents to learn from unsuccessful attempts and progressively improve.
@@
Process-based Reward In contrast to outcome-based rewards, which provide a single, delayed signal, the process-based reward paradigm offers more precise and granular feedback by evaluating each step in an agent’s trajectory. Process-supervised reward models (PRMs) have been demonstrated to be significantly more reliable than outcome-supervised reward models (ORMs), particularly in domains requiring complex reasoning like solving math problems lightman2023let . However, obtaining such fine-grained step-level feedback traditionally requires extensive human annotations, which are both time-consuming and expensive to scale. To address this annotation bottleneck, Math-Shepherd wang2023math proposes an automatic process annotation framework that utilizes Monte Carlo Tree Search (MCTS) to gather step-wise supervision by assessing each step’s potential to derive the correct final answer. Similarly, AlphaMath chen2024alphamath trains a value model to evaluate the step correctness in solution paths and updates both the policy and value model through exploration and exploitation within an MCTS framework. By leveraging process-based rewards, agents can improve their capabilities in a progressive, step-by-step manner. rStar-Math guan2025rstar and AgentPRM choudhury2025process both propose methods to iteratively evolve the policy and the process reward model, generating progressively higher-quality reasoning paths without manual labels. Agent Q putta2024agentq integrates a step-wise verification mechanism into its MCTS process to collect high-quality trajectories, which are then used to iteratively refine the policy via DPO training.
@@
Hybrid Reward The hybrid methods aim to provide more comprehensive learning signals by incorporating both the clarity of final task success (outcome-based) and the granular guidance of intermediate steps (process-based). These methods overcome the sparsity of outcome-only signals while grounding the agent’s step-by-step reasoning in the ultimate task goal. For example, GiGPO feng2025group addresses the instability of training long-horizon agents by introducing a dual-level reward mechanism. It provides an episode-level reward based on the final success of entire trajectories, while simultaneously assigning a localized, step-level reward for intermediate actions. This dual signal provides both a high-level directional goal and low-level corrective guidance. Similarly, SPA-RL wang2025spa proposes a reward decomposition method that bridges the gap between sparse outcome signals and dense process feedback. It attributes incremental progress to each step within multi-step trajectories based on the final task completion, effectively distributing the outcome-based reward across the process steps. This approach creates dense intermediate progress rewards that enhance reinforcement learning effectiveness while maintaining alignment with the ultimate task objectives.
@@ 6.1 General Domain Evolution
The first category, general domain evolution, refers to self-evolving agents designed for general-purpose applications, particularly as versatile digital assistants. These agents progressively enhance their capabilities to address a broad spectrum of user queries, especially in dynamic and diverse digital environments. Technically speaking, these general assistant agent enhance their abilities primarily via three mechanisms: memory optimization, curriculum-driven training, and model-agent co-evolution. These mechanisms collectively enable the agents to continuously adapt and effectively respond to increasingly complex user demands.
@@
The most common mechanism facilitating agent evolution is the memory mechanism, wherein agents summarize historical success/failure experiences (wang2023voyager, ; zhang2024offline, ) into memory representations (zhang2024survey, ), anticipating that these distilled experiences will be beneficial when addressing previously unseen tasks. For instance, Mobile-Agent-E (wang2025mobile, ) employs a long-term memory structure consisting of "Tips," which provide general guidelines, and "Shortcuts," representing reusable action sequences derived from past experiences. This self-evolutionary module supports the continuous enhancement of performance on complex smartphone tasks. Another typical example is MobileSteward (liu2025mobilesteward, ), which coordinates multiple app-specific Agents under a central Agent, with specialized modules for task scheduling, execution, and evaluation. It also incorporates a memory-based self-evolution mechanism that summarizes successful executions to improve future cross-app instruction handling. Meanwhile, Generative Agents (park2023generative, ) store episodic memories of their experiences, synthesize higher-level reflections, and condition future planning on this self-reflection. In these examples, memory serves as the foundation that enables agents to internalize past experiences, abstract high-level patterns, and refine their future behavior.
@@
Model-Agent Co-Evolution.
Another line of work is to perform Model-Agent Co-evolution for LLM agents. UI-Genie (xiao2025ui, ) constructs a specialized image-text reward model that scores trajectories at both step and task levels. It jointly fine-tunes the agent and reward model using synthetic trajectories—generated by controlled corruption and hard-negative mining—across multiple generations. WebEvolver (fang2025webevolver, ) introduces a co-evolving world model LLM that simulates web environments. It generates synthetic training data by predicting next observations and enables look-ahead reasoning during inference, which greatly improves real-web task success. Absolute Zero (zhao2025absolute, ) co-evolves a reasoning agent and its internal self-reward model through reinforced self-play. By adversarially generating increasingly challenging reasoning problems and optimizing the agent using internal self-certainty as a reward signal, the framework simultaneously updates both the agent’s policy and the self-rewarding mechanism. Together, these methods demonstrate the effectiveness of co-evolving agents and auxiliary models (e.g., reward or world models) to achieve more robust, generalizable, and scalable learning in LLM agentic systems.
@@
Curriculum-driven training also serves as a critical mechanism for building a self-evolving general assistant. For example, WebRL (qi2024webrl, ) uses a self-evolving curriculum: when an agent fails, similar but manageable tasks are automatically generated. Coupled with a learned reward model and adaptive policy updates, this yields a success rate uplift on WebArena benchmarks. Voyager (wang2023voyager, ) similarly leverages an automatic, bottom-up curriculum in Minecraft, where GPT-4 proposes appropriate next tasks based on agent progress, building a growing code-based skill library through iterative prompting and environmental feedback. These approaches highlight how curriculum learning enables agents to autonomously expand their capabilities through iterative task adaptation.

### v4

@@ 5.3.2 Multi-Agent Evolution
System Architecture Evolution. This paradigm focuses on evolving the structural and coordination aspects of multi-agent systems, including team composition, orchestration strategies, and workflow optimization. EvoMAC (Hu et al., 2024d) introduces a framework that mimics neural network training for multi-agent systems, implementing "textual backpropagation" where compilation errors and test failures serve as loss signals to drive iterative modifications of agent team composition and individual prompts. A specialized "updating team" analyzes textual feedback to identify problematic agents and generate modification instructions, effectively implementing gradient-based optimization in the space of agent configurations rather than model parameters. The FELA framework (Wang et al., 2025e) applies this concept to a practical industrial problem, using a multi-agent system with specialized "Idea," "Code," and "Critic" agents that collaboratively evolve to generate high-performing features from complex data, guided by principles from both reinforcement learning and genetic algorithms. Puppeteer (Dang et al., 2025) takes a different approach by focusing on coordination strategy evolution rather than team composition changes. The system employs a centralized orchestrator that evolves its decision policy through reinforcement learning, dynamically selecting which agents to activate at each step while balancing task performance with computational cost. This "puppeteer-puppet" paradigm demonstrates how architectural evolution can occur at the coordination level, discovering efficient collaboration patterns and emergent behaviors such as tighter coordination among core agents and sophisticated cyclic interaction patterns. Agent0 (Xia et al., 2025b) introduces a framework that evolves agents from zero data via a co-evolutionary loop between a curriculum agent and an executor agent. The curriculum agent is trained to generate frontier tasks that challenge the executor. Then, the improved tool-use capabilities of the executor in turn drive the creation of a more complex, tool-aware curriculum, establishing a virtuous cycle of self-improvement.
@@
Knowledge-Based Evolution. This paradigm emphasizes evolving the collective knowledge and experience of multi-agent teams through memory accumulation and case-based learning, primarily operating through in-context learning or in-context-like adaptation rather than parameter updates. MDTeamGPT (Chen et al., 2025e) establishes the foundation for this approach through a dual knowledge base system, implementing CorrectKB for storing successful cases and ChainKB for capturing failure reflections, enabling the system to learn from both successes and mistakes through structured case retrieval and reasoning enhancement. Extending this medical consultation framework, MedAgentSim (Almansoori et al., 2025a) demonstrates how such knowledge-based evolution can be applied to real-world diagnostic scenarios, accumulating experience from patient interactions and using retrieval-augmented generation to improve consultation quality over time. PiFlow (Pu et al., 2025) applies this paradigm to scientific discovery, maintaining a trajectory of principle-outcome pairs and using them to steer hypothesis generation through information-theoretical optimization.
In summary, population-based and self-play evolution enhance diversity and open-ended discovery, yet typically incur higher computational cost and lower interpretability compared with single-agent paradigms.
@@ 5.4 Cross-cutting Evolutionary Dimensions
After outlining the three core evolutionary paradigms, we now analyze their cross-cutting dimensions—revealing how different design choices balance feedback type, data source, and learning stability.
Figure 7: Illustration of the cross-cutting evolutionary dimensions underlying agent self-evolution. Learning Paradigm: in offline learning, data are pre-collected and used for filtering and fine-tuning, while in online learning, agents continuously interact with their environments for real-time adaptation. Policy Consistency: on-policy evolution updates policies based on the agent’s own trajectories, whereas off-policy evolution relies on replay buffers, human demonstrations, or experiences from other agents. Reward Granularity: feedback can be process-based (step-level rewards), outcome-based (final-result rewards), or a hybrid of both. Together, these three orthogonal axes define how agents generate data, adapt their policies, and receive feedback across reward-based, imitation-based, and population-based evolution paradigms.
@@ 5.4.1 Online and Offline Learning
Offline Learning In the offline learning paradigm, the learning phase is decoupled from live task execution. The offline process typically involves cycles of offline data generation, filtering, and model fine-tuning, focusing on building a powerful and generalist foundational model before deployment. A primary strategy in this domain is LLM bootstrapping, where a model enhances its own capabilities using its self-generated content. For example, Self-Instruct(Wang et al., 2022) shows how a language model can bootstrap its own instruction-following ability by generating new instructions, paired with its own responses, creating a synthetic dataset for fine-tuning. Building on this, WizardLM(Xu et al., 2024a) demonstrates how to progressively evolve the complexity of these self-generated instructions, pushing the model’s capabilities on more challenging tasks. Although these methods primarily focus on broad capability expansion via synthetic heuristics, acting as a bootstrapping phase, they lay the necessary groundwork for closed-loop, experience-driven evolution defined in our framework. In the context of GUI and Web agents, offline learning often involves leveraging pre-collected high-quality trajectories for supervised fine-tuning (SFT). OS-Genesis(Sun et al., 2024b) introduced a reverse task synthesis method for automatic trajectory creation. Similarly, UI-Genie(Xiao et al., 2025a) employs a unified reward model for trajectory evaluation and a self-improving loop to generate high-quality trajectories iteratively. Both approaches focus on curating a rich SFT dataset to enhance the agent’s capabilities to solve complex tasks. Beyond SFT, offline methods also incorporate reinforcement learning performed on a static dataset of agent-environment interactions. For example, GUI-R1(Luo et al., 2025b) and InfiGUI-R1(Liu et al., 2025e) utilize rule-based rewards and apply R1-style(Guo et al., 2025b) training on offline GUI datasets.
@@
Online Learning In contrast, online learning enables an agent to learn and adapt continuously while it interacts with a live or simulated environment. Feedback from each action is used to update the agent’s policy, plan, or knowledge base in real-time. This allows for greater adaptability to dynamic or unseen situations. Some agents evolve online not by updating their model weights, but by refining their plans and skill libraries on the fly. For example, Voyager(Wang et al., 2023a) presents an LLM-powered agent that learns to play Minecraft by continuously exploring, generating its own curriculum of tasks, and building a persistent skill library from direct experience. AdaPlanner(Sun et al., 2023) focuses on adapting its plan within a task; it generates an initial plan, receives feedback from the environment, and refines the plan online. Similarly, SwiftSage(Lin et al., 2023) operates with a fast-and-slow thinking process, where it can reflect on failures of its fast, intuitive mode and switch to a more deliberate, tool-using slow mode, adapting its strategy online based on task difficulty. Reinforcement Learning serves as a fundamental mechanism for online learning, enabling agents to learn from environmental reward signals. DigiRL(Bai et al., 2024) demonstrates how to train device-control agents in the wild using autonomous RL, while DistRL(Wang et al., 2024g) proposes an asynchronous distributed framework to make such on-device training feasible. MobileGUI-RL(Shi et al., 2025b) addresses the specific challenges of training GUI agents in online mobile environments by introducing a synthetic task generation pipeline combined with group relative policy optimization (GRPO) through trajectory-aware rewards.
@@ 5.4.2 On-policy and Off-policy Learning
On-policy Learning. On-policy approaches require agents to learn exclusively from experiences generated by their current policy, ensuring policy consistency but often at the cost of sample efficiency. Reflexion (Shinn et al., 2023) exemplifies this approach through its iterative self-reflection mechanism. The agent generates responses using its current policy, receives feedback on failures, and immediately incorporates this feedback to update its reasoning process for the next iteration. GRPO (Shao et al., 2024b) and DAPO (Yu et al., 2025a) continue this path and show the effectiveness of multiple rollouts. The agent always learns from its current behavior, maintaining strict policy consistency. In agent settings, on-policy methods provide excellent learning stability and avoid distribution mismatch issues that plague off-policy methods. However, they suffer from low sample efficiency, as each policy update requires fresh data collection, making them computationally expensive for complex multi-step reasoning or tool use scenarios where generating high-quality trajectories is costly.
Off-policy Learning. Off-policy approaches allow agents to learn from experiences generated by different policies, including previous versions, other agents, or human demonstrations, significantly improving sample efficiency at the cost of potential distribution mismatch. Yuan et al. (2024d) demonstrates a sophisticated off-policy approach where model $M_{t+1}$ learns from preference data generated by the previous version $M_{t}$. The system handles distribution shift through DPO’s built-in KL divergence constraint with the reference policy, preventing the new policy from deviating too far from the data-generating policy. Yuan et al. (2023) showcases another powerful off-policy paradigm by learning from diverse response sources—including other models, humans, and different sampling strategies—through ranking-based supervision. The method elegantly sidesteps distribution shift by treating alignment as a ranking problem rather than requiring policy consistency. Zhao et al. (2025b) illustrates off-policy learning in multi-agent settings, where agents learn from an "experience library" containing successful interaction trajectories generated by previous policy versions, enabling efficient reuse of expensive multi-agent coordination data. In agent settings, off-policy methods excel in sample efficiency, allowing agents to leverage historical data, expert demonstrations, and cross-agent learning. They are particularly valuable for multi-step reasoning where successful trajectories are rare and expensive to generate, and for tool use scenarios where agents can learn from diverse execution examples without repeated environmental interaction. However, they face challenges with distribution shift, reward hacking (where agents exploit inconsistencies between training and deployment policies), and the need for careful regularization to maintain training stability.
@@ 5.4.3 Reward Granularity
Outcome-based Reward Outcome-based Reward is a feedback mechanism that evaluates an agent based on the successful completion of predefined tasks. This reward is determined solely by the final state of the agent’s trajectory, regardless of the intermediate steps. A central challenge, particularly in dynamic environments like web or GUI navigation, is to effectively learn from both successful trajectories and the much more frequent failure trajectories. To address this, Direct Preference Optimization (DPO)(Rafailov et al., 2023) is designed to directly maximize the likelihood of preferred responses while minimizing the KL-divergence with the reference policy. Similarly, RRHF(Yuan et al., 2023) employs a ranking loss approach that aligns model probabilities of multiple responses with human preferences by ranking response probabilities without requiring auxiliary value models. Moreover, several works have developed specialized frameworks for agent self-evolution that are built upon outcome-based rewards. A straightforward approach is rejection sampling finetuning, as used in AutoWebGLM(Lai et al., 2024). This method employs a pre-designed reward model to evaluate trajectory outcomes, identify the successful trajectories, and update the model with this high-quality data. DigiRL(Bai et al., 2024) models the GUI navigation task as a Markov Decision Process (MDP) and obtains a final, sparse reward at the end of an episode using a VLM-based evaluator. WebRL(Qi et al., 2024) develops a robust outcome-supervised reward model (ORM) to address the feedback sparsity inherent in dynamic web environments. The ORM evaluates task success within a self-evolving curriculum framework, enabling agents to learn from unsuccessful attempts and progressively improve.
@@
Process-based Reward In contrast to outcome-based rewards, which provide a single, delayed signal, the process-based reward paradigm offers more precise and granular feedback by evaluating each step in an agent’s trajectory. Process-supervised reward models (PRMs) have been demonstrated to be significantly more reliable than outcome-supervised reward models (ORMs), particularly in domains requiring complex reasoning like solving math problems(Lightman et al., 2023). However, obtaining such fine-grained step-level feedback traditionally requires extensive human annotations, which are both time-consuming and expensive to scale. To address this annotation bottleneck, Math-Shepherd(Wang et al., 2023b) proposes an automatic process annotation framework that utilizes Monte Carlo Tree Search (MCTS) to gather step-wise supervision by assessing each step’s potential to derive the correct final answer. Similarly, AlphaMath(Chen et al., 2024a) trains a value model to evaluate the step correctness in solution paths and updates both the policy and value model through exploration and exploitation within an MCTS framework. By leveraging process-based rewards, agents can improve their capabilities in a progressive, step-by-step manner. rStar-Math(Guan et al., 2025) and AgentPRM(Choudhury, 2025) both propose methods to iteratively evolve the policy and the process reward model, generating progressively higher-quality reasoning paths without manual labels. Agent Q(Putta et al., 2024) integrates a step-wise verification mechanism into its MCTS process to collect high-quality trajectories, which are then used to iteratively refine the policy via DPO training.
@@
Hybrid Reward The hybrid methods aim to provide more comprehensive learning signals by incorporating both the clarity of final task success (outcome-based) and the granular guidance of intermediate steps (process-based). These methods overcome the sparsity of outcome-only signals while grounding the agent’s step-by-step reasoning in the ultimate task goal. For example, GiGPO(Feng et al., 2025a) addresses the instability of training long-horizon agents by introducing a dual-level reward mechanism. It provides an episode-level reward based on the final success of entire trajectories, while simultaneously assigning a localized, step-level reward for intermediate actions. This dual signal provides both a high-level directional goal and low-level corrective guidance. Similarly, SPA-RL(Wang et al., 2025d) proposes a reward decomposition method that bridges the gap between sparse outcome signals and dense process feedback. It attributes incremental progress to each step within multi-step trajectories based on the final task completion, effectively distributing the outcome-based reward across the process steps. This approach creates dense intermediate progress rewards that enhance reinforcement learning effectiveness while maintaining alignment with the ultimate task objectives.
@@ 6.1 General Domain Evolution
The first category, general domain evolution, refers to self-evolving agents designed for general-purpose applications, particularly as versatile digital assistants. These agents progressively enhance their capabilities to address a broad spectrum of user queries, especially in dynamic and diverse digital environments. Technically speaking, these general assistant agent enhance their abilities primarily via three mechanisms: memory optimization, curriculum-driven training, and model-agent co‑evolution. These mechanisms collectively enable the agents to continuously adapt and effectively respond to increasingly complex user demands.
@@
The most common mechanism facilitating agent evolution is the memory mechanism, wherein agents summarize historical success/failure experiences (Wang et al., 2023a; Zhang et al., 2024f) into memory representations (Zhang et al., 2024g), anticipating that these distilled experiences will be beneficial when addressing previously unseen tasks. For instance, Mobile-Agent-E (Wang et al., 2025n) employs a long-term memory structure consisting of "Tips," which provide general guidelines, and "Shortcuts," representing reusable action sequences derived from past experiences. This self-evolutionary module supports the continuous enhancement of performance on complex smartphone tasks. Another typical example is MobileSteward (Liu et al., 2025f), which coordinates multiple app-specific Agents under a central Agent, with specialized modules for task scheduling, execution, and evaluation. It also incorporates a memory-based self-evolution mechanism that summarizes successful executions to improve future cross-app instruction handling. Meanwhile, Generative Agents (Park et al., 2023) store episodic memories of their experiences, synthesize higher-level reflections, and condition future planning on this self-reflection. In these examples, memory serves as the foundation that enables agents to internalize past experiences, abstract high-level patterns, and refine their future behavior.
@@
Model-Agent Co‑Evolution.
Another line of work is to perform Model-Agent Co‑evolution for LLM agents. UI‑Genie (Xiao et al., 2025a) constructs a specialized image-text reward model that scores trajectories at both step and task levels. It jointly fine-tunes the agent and reward model using synthetic trajectories—generated by controlled corruption and hard-negative mining—across multiple generations. WebEvolver (Fang et al., 2025b) introduces a co-evolving world model LLM that simulates web environments. It generates synthetic training data by predicting next observations and enables look-ahead reasoning during inference, which greatly improves real-web task success. Absolute Zero (Zhao et al., 2025a) co‑evolves a reasoning agent and its internal self‑reward model through reinforced self‑play. By adversarially generating increasingly challenging reasoning problems and optimizing the agent using internal self‑certainty as a reward signal, the framework simultaneously updates both the agent’s policy and the self-rewarding mechanism. Together, these methods demonstrate the effectiveness of co-evolving agents and auxiliary models (e.g., reward or world models) to achieve more robust, generalizable, and scalable learning in LLM agentic systems.
@@
Curriculum-driven training also serves as a critical mechanism for building a self-evolving general assistant. For example, WebRL (Qi et al., 2024) uses a self-evolving curriculum: when an agent fails, similar but manageable tasks are automatically generated. Coupled with a learned reward model and adaptive policy updates, this yields a success rate uplift on WebArena benchmarks. Voyager (Wang et al., 2023a) similarly leverages an automatic, bottom‑up curriculum in Minecraft, where GPT‑4 proposes appropriate next tasks based on agent progress, building a growing code-based skill library through iterative prompting and environmental feedback. These approaches highlight how curriculum learning enables agents to autonomously expand their capabilities through iterative task adaptation.